If m,n are roots of equation x2+mn=(m+1)x 
Then n equals to 

Grade:12

2 Answers

Arun
5 years ago
Dear Priya
 
sum = m + n
 
now from equation
 
x^2 – (m + 1)x + mn = 0
 
hence sum = m + 1 = m+ n
 
so 
n = 1
 
Regards
Arun
Deepak Kumar Shringi
5 years ago
x^2-(m+1)x+mn=0
so m+n=m+1
n=1